# Guardrailed Uplift Targeting: A Causal Optimization Playbook for Marketing Strategy

### v1

@@ Abstract
This paper introduces a marketing decision framework that converts heterogeneous-treatment uplift into constrained targeting strategies to maximize revenue and retention while honoring business guardrails. The approach estimates Conditional Average Treatment Effects (CATE) with uplift learners and then solves a constrained allocation to decide who to target and which offer to deploy under limits such as budget or acceptable sales deterioration. Applied to retention messaging, event rewards, and spend-threshold assignment, the framework consistently outperforms propensity and static baselines in offline evaluations using uplift AUC, Inverse Propensity Scoring (IPS), and Self-Normalized IPS (SNIPS). A production-scale online A/B test further validates strategic lift on revenue and completion while preserving customer-experience constraints. The result is a reusable playbook for marketers to operationalize causal targeting at scale, set guardrails, and align campaigns with strategic KPIs.
@@ I Introduction
The emergence of causal machine learning has enabled more sophisticated approaches to treatment effect estimation. Uplift modeling techniques, including meta-learners and forest-based estimators, provide methods for estimating individual treatment effects [22, 28]. However, translating these estimates into actionable targeting strategies that respect business constraints remains challenging.
@@
The remainder of this paper is organized as follows: Section 2 reviews related work in uplift modeling and constrained optimization. Section 3 details our methodological framework. Section 4 presents experimental applications and results. Section 5 discusses implementation considerations, and Section 6 concludes with implications for marketing practice.
@@ II Related Work
Our framework for guardrailed uplift targeting intersects with several streams of research in machine learning, data systems, and optimization. While traditional uplift modeling has focused primarily on treatment effect estimation [19], our approach extends this by integrating business constraints into the optimization process.
@@ III-B Uplift Estimation Stage
where $Y_{i}(1)$ and $Y_{i}(0)$ represent potential outcomes under treatment and control, respectively.
@@
  • Causal Forest: An adaptation of random forests for treatment effect estimation that handles high-dimensional covariates effectively [28].
@@ III-D Evaluation Methodology
IV Experimental Applications and Results
IV-A Customer Retention Optimization
IV-A1 Business Context and Challenge
@@ IV-A3 Results and Analysis
IV-B Event Revenue Maximization
IV-B1 Business Context and Challenge
@@
In our application, a platform offered two reward levels ($P_{1}$ and $P_{2}$ with $P_{1}<P_{2}$) for purchases exceeding a threshold amount during promotional events. The historical approach randomly assigned reward levels, ignoring customer heterogeneity in response to incentives.
IV-B2 Methodology and Constraints
@@
IV-B3 Results and Business Impact
@@
Our optimized policy achieved a 0.317% lift in revenue compared to the completion score baseline while reducing the efficiency metric (e%iS) by 11.879%. The policy also improved completion rates by 1.02%, indicating better customer experience alongside financial improvements.
@@
IV-C Spend Threshold Optimization
IV-C1 Business Context
@@ IV-C2 Experimental Approach
IV-C3 Online Validation Results
@@
V Implementation Considerations
V-A Data Requirements and Experimental Design
@@
V-B Model Selection and Validation
@@
V-C Constraint Specification and Sensitivity
@@ V-D Computational Scalability
VI Conclusion and Future Directions
VI-A Summary of Contributions
This paper presents a comprehensive framework for guardrailed uplift targeting that combines causal inference with constrained optimization. Our approach addresses key limitations of traditional targeting methods by incorporating heterogeneous treatment effects and business constraints into a unified decision-making process.
The experimental results demonstrate consistent improvements over baseline approaches across multiple business scenarios. In retention optimization, our framework achieved 2.35-3.905% lifts while reducing targeting scope. In revenue maximization, it improved financial metrics while maintaining customer experience. The online validation confirmed these benefits in a production environment, with statistically significant improvements in both revenue and engagement metrics.
VI-B Practical Implications
The framework provides marketers with a systematic approach to campaign optimization that balances multiple business objectives. By explicitly incorporating constraints, it enables more realistic and implementable targeting strategies that respect operational limitations and strategic priorities.
The methodology supports personalization at scale, moving beyond one-size-fits-all approaches while maintaining control over business outcomes. The reusable playbook structure facilitates adoption across different marketing domains, from retention and acquisition to promotional optimization.
VI-C Limitations and Extensions
While our framework demonstrates strong performance across applications, several limitations warrant attention. The unconfoundedness assumption, while common in causal inference, may be violated in practice. Future work could incorporate methods for handling unobserved confounding, such as sensitivity analysis [18] or proximal causal inference [27].
The current implementation focuses on binary and discrete treatments, but many marketing interventions involve continuous parameters (e.g., discount percentage). Extending the framework to continuous treatments using methods like double/debiased machine learning [17] represents a promising direction.
Additional research could explore dynamic targeting policies that adapt to changing customer behavior and market conditions. Reinforcement learning approaches integrated with causal inference could provide more responsive optimization while maintaining the guardrails essential for business applications.
VI-D Concluding Remarks
Guardrailed uplift targeting represents a significant advance in marketing optimization, bridging the gap between causal inference theory and practical business decision-making. By combining sophisticated uplift modeling with flexible constraint handling, the framework enables more effective, efficient, and responsible marketing personalization.
As e-commerce continues to evolve, approaches that balance personalization with business constraints will become increasingly valuable. Our framework provides a foundation for this evolution, supporting data-driven marketing strategies that maximize value for both businesses and customers.

### v2

@@ Abstract
This paper introduces a marketing decision framework that optimizes customer targeting by integrating heterogeneous treatment effect estimation with explicit business guardrails. The objective is to maximize revenue and retention while adhering to constraints such as budget, revenue protection, and customer experience. The framework first estimates Conditional Average Treatment Effects (CATE) using uplift learners, then solves a constrained allocation problem to decide whom to target and which offer to deploy. It supports decisions in retention messaging, event rewards, and spend-threshold assignment. Validated through offline simulations and online A/B tests, the approach consistently outperforms propensity and static baselines, offering a reusable playbook for causal targeting at scale.
@@ I Introduction
The emergence of causal machine learning has enabled more sophisticated approaches to treatment effect estimation. Uplift modeling techniques, including meta-learners and forest-based estimators, provide methods for estimating individual treatment effects [22, 28]. Recent work has demonstrated the value of combining uplift modeling with optimization frameworks in e-commerce settings, including knapsack formulations for promotion personalization [13] and retrospective uplift modeling within ROI constraints [20]. Extensions to multiple treatments with explicit cost optimization have further advanced practical applicability [29]. However, translating these estimates into actionable targeting strategies that systematically respect diverse business constraints remains challenging.
@@
The remainder of this paper is organized as follows: Section 2 reviews related work in uplift modeling and constrained optimization. Section 3 details our methodological framework. Section 4 describes the simulation setup for offline validation. Section 5 presents experimental applications and results. Section 6 discusses implementation considerations, and Section 7 concludes with implications for marketing practice.
I-A Contributions
Our main contributions are:
  • A two-stage guardrailed uplift targeting framework that combines CATE estimation with constrained optimization for practical marketing decisions.
  • A detailed simulation setup for offline validation, including data generation, treatment assignment, response functions, and guardrail constraints.
  • Empirical validation across three business scenarios (retention, event revenue, spend thresholds) showing consistent improvement over baselines.
  • Implementation guidelines covering data requirements, model selection, constraint specification, and computational scalability.
  • Publicly available code and a reusable playbook for marketers to operationalize causal targeting with guardrails.
@@ II Related Work
Our framework for guardrailed uplift targeting intersects with several streams of research in machine learning, data systems, and optimization. Foundational work in uplift modeling established core methodologies for treatment effect estimation, including comprehensive surveys of the field [21] and the development of specialized decision tree algorithms for single and multiple treatment scenarios [25]. While traditional uplift modeling has focused primarily on treatment effect estimation [19], our approach extends this by integrating business constraints into the optimization process.
@@ III-B Uplift Estimation Stage
where $Y_{i}(1)$ and $Y_{i}(0)$ represent potential outcomes under treatment and control, respectively. This notation follows the potential outcomes framework formalized by Rubin [24], which provides the theoretical foundation for causal inference from experimental and observational data.
@@
  • Causal Forest: An adaptation of random forests for treatment effect estimation that handles high-dimensional covariates effectively [28]. Related approaches include orthogonal random forests, which achieve improved convergence properties through Neyman-orthogonal score functions [23].
@@ III-D Evaluation Methodology
IV SIMULATION SETUP
IV-A Data Generation and Treatment Assignment
We simulate customer data with $N=100,000$ observations and $p=20$ covariates, drawn from a multivariate normal distribution with correlation $\rho=0.3$. A binary treatment $T_{i}\in\{0,1\}$ is assigned via a randomized trial with propensity $P(T_{i}=1)=0.5$. The response function follows:
$Y_{i}=\alpha+\beta X_{i}+\tau(X_{i})T_{i}+\epsilon_{i},$
where $\tau(X_{i})$ is the heterogeneous treatment effect modeled as $\tau(X_{i})=\gamma_{0}+\gamma_{1}X_{i1}+\gamma_{2}X_{i2}^{2}$, and $\epsilon_{i}\sim\mathcal{N}(0,\sigma^{2})$.
IV-B Guardrail Constraints
We impose three business guardrails:
  1. 1.
    Budget constraint: At most $B=10\%$ of customers may receive treatment.
  2. 2.
    Revenue protection: Total revenue under the policy must not drop by more than $\delta=2\%$ compared to treating all.
  3. 3.
    Fairness constraint: Demographic parity gap $\leq 0.05$ across two protected groups.
IV-C Evaluation Protocol
We evaluate using:
  • Uplift AUC: Model discrimination ability.
  • ROI: Return on investment relative to cost.
  • Constraint violation rate: Percentage of simulations violating any guardrail.
  • Baselines: Propensity score targeting, random assignment, and treat-all.
  • Sensitivity checks: Varying noise levels $\sigma^{2}$ and treatment effect heterogeneity.
V Experimental Applications and Results
V-A Customer Retention Optimization
V-A1 Business Context and Challenge
@@ V-A3 Results and Analysis
V-B Event Revenue Maximization
V-B1 Business Context and Challenge
@@
In our application, a platform offered two reward levels ($P_{1}$ and $P_{2}$ with $P_{1}<P_{2}$) for purchases exceeding a threshold amount during promotional events. The historical approach randomly assigned reward levels, ignoring customer heterogeneity in response to incentives. Prior work has demonstrated that personalized reward assignment using uplift modeling can significantly improve promotional efficiency [13], [20], particularly when combined with cost optimization [29].
V-B2 Methodology and Constraints
@@
V-B3 Results and Business Impact
@@
Our optimized policy achieved a 0.317% lift in revenue compared to the completion score baseline while reducing the efficiency metric which is expected Incentive Spend as a percentage of Sales (e%iS) by 11.879%. The policy also improved completion rates by 1.02%, indicating better customer experience alongside financial improvements.
@@
V-C Spend Threshold Optimization
V-C1 Business Context
@@ V-C2 Experimental Approach
V-C3 Online Validation Results
@@
VI Implementation Considerations
VI-A Data Requirements and Experimental Design
@@
VI-B Model Selection and Validation
@@
VI-C Constraint Specification and Sensitivity
@@ VI-D Computational Scalability
VII CONCLUSION AND FUTURE WORK
VII-A Summary
We introduced a guardrailed uplift targeting framework that combines causal uplift modeling with constrained optimization. The method outperforms traditional baselines across retention, revenue, and threshold optimization tasks while respecting business guardrails. Online A/B tests confirm statistically significant lifts in revenue and engagement.
VII-B Future Directions
  • Deployment considerations: Real-world A/B testing protocols for dynamic policy updates.
  • Robustness to unobserved confounding: Integrating sensitivity analysis to assess omitted variable bias [18] or proximal causal inference methods that leverage proxy variables [27].
  • Multi-period targeting: Extending to sequential decision-making with reinforcement learning.
  • Continuous treatments: Adapting the framework for continuous incentives (e.g., discount percentages) using doubly robust methods.